OpenStudy (anonymous):

Are the following lines parallel, perpendicular, or neither? y = -7x + 12 - 7y = x + 56 Select one: a. Neither b. Perpendicular c. Parallel

OpenStudy (anonymous):

a. Neither

OpenStudy (anonymous):

they are not besides each other and when they meet they do not form a 90 degree angle

OpenStudy (shamim):

u hv to compare ur equation of line with standard equation

OpenStudy (shamim):

standard equation of line is y=mx+c

OpenStudy (shamim):

so the slope of ur first equation of line is m1=-7

OpenStudy (shamim):

and the slope of ur second equation of line is m2=-1/7

OpenStudy (shamim):

2 lines will b parallel if m1=m2

OpenStudy (shamim):

2 lines will b perpendicular with each other if m1*m2=-1

OpenStudy (shamim):

a